OpenStudy (anonymous):

If a shark's tooth is shaped like a triangle the height of the is 3cm longer than the base .The area is 14cm2 how do i find the height and base?

OpenStudy (anonymous):

what's the formula for the area of a triangle?

OpenStudy (anonymous):

14cm squared

OpenStudy (anonymous):

no, what is the formula to figure out a triangle's area from its base and height?

OpenStudy (anonymous):

this was a word question so i didn't see a formula

OpenStudy (anonymous):

no, i'm asking you if you know the formula to find a triangle's area from its base and height!

OpenStudy (anonymous):

you know, like area of a square is side x side? what's area of a triangle?

OpenStudy (anonymous):

no it's to find the sharks teeth length that are shaped like a triangle

OpenStudy (anonymous):

I KNOW! WHAT IS THE FORMULA FOR THE AREA OF A TRIANGLE?!

OpenStudy (anonymous):

I'm trying to helpfully lead you to the solution, but you're not really cooperating.

OpenStudy (anonymous):

14cm2*14cm2=

OpenStudy (anonymous):

OK, I'll give it to you. The area of a triangle is 1/2 base x height.

OpenStudy (anonymous):

and you know height = base + 3, and area is 14.

OpenStudy (anonymous):

so is it 14*3=

OpenStudy (anonymous):

no. Area = 1/2 base * height. So it's 14 = 1/2(base)(base + 3)

OpenStudy (anonymous):

because 14 is the area, base is what we're solving for, and height is base + 3.

OpenStudy (anonymous):

1/2*14cm2*3=

OpenStudy (anonymous):

no. you multiply out the earlier equation to get 28 = base2 + 3base. And then base2 + 3base - 28 = 0. Then you factor into (base + 7)(base - 4) = 0.

OpenStudy (anonymous):

okay I think I got it now .

OpenStudy (anonymous):

OK, good. :-)

OpenStudy (anonymous):

Thank you